OpenStudy (anonymous):

jodie bicycles 7 km/h faster than carlos. in the same time it takes carlos to bicycle 57km, jodie can bicycle 78 km. how fast does each bicyclist travel?

OpenStudy (anonymous):

let x = carlos's speed jodie: x + 7 x * t = 57 (x + 7) * t = 78 solve for t and equate t = 57/x t = 78 / (x + 7) \[\frac{57}{x} = \frac{78}{x + 7}\] 57(x + 7) = 78x 57x + 399 = 78x 21x = 399 x = 19 carlos bikes at 19 km/hr jodie bikes at 26 km/hr